OpenStudy (hitaro9):

Assume that abs[f '(x)] ≤ M for a ≤ x ≤ b. One can show by the Mean Value Theorem that abs[f(b) − f(a)] ≤ M (b − a), and thus f(a) − M(b − a) ≤ f(b) ≤ f(a) + M(b − a). Using this fact, determine a lower bound and an upper bound for sqrt101 (Hint: Let f (x) = sqrt(x), a = 100, and b = 101, and then find bounds for f(101) = sqrt(101). First, find a reasonable value for M: M = ______

OpenStudy (hitaro9):

No idea what M even is. We didn't talk a bout this in class and its not in the textbook

OpenStudy (hitaro9):

@wio please help <.>

OpenStudy (hitaro9):

So like, I would think that M could be any almost any positive number for the equation to be true, but I don't know what is considered "reasonable"

OpenStudy (anonymous):

\[ |f'(x)|<M \]Using mean value theorem on this gives us \[ \left|\frac{f(b)-f(a)}{b-a}\right|<M \]

OpenStudy (anonymous):

And \[ \frac{|f(b)-f(a)|}{|b-a|} \]Now since \(a\leq b\) then \(0\leq b-a\) so \(b-a\) is zero or positive meaning \(|b-a|=b-a\)

OpenStudy (anonymous):

\[ \frac{|f(b)-f(a)|}{b-a}<M \]And \[ |f(b)-f(a)|<M(b-a) \]

OpenStudy (anonymous):

\(M\) is just some number, a bound for the derivative

OpenStudy (hitaro9):

So sqrt 101- sqrt 100? And then just plug that in?

OpenStudy (hitaro9):

I thought that too satellite, but I plugged in 1, and it said that that was wrong.

OpenStudy (anonymous):

We have two equations: \[ f(b)-f(a)<M(b-a) \]And \[ f(a)-f(b)<M(b-a) \]

OpenStudy (anonymous):

no it is not 1 you are asked for a bound for \(f'(x)=\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

since \(x>100\) i guess you can use \(\frac{1}{2\sqrt{100}}=\frac{1}{20}\)

OpenStudy (anonymous):

that is just a guess, mind you, i would not bet on it

OpenStudy (anonymous):

\[ f'(x) \]Is a monotonic decreasing function, so it is highest at the beginning of any interval.

OpenStudy (anonymous):

So letting \(M=f'(100)\) is fine here.

OpenStudy (anonymous):

And \(b-a=1\)

OpenStudy (hitaro9):

Okay, I see why 1/20 satisfies the f(b)-f(a) < M part, but why can it not be greater than that?

OpenStudy (anonymous):

No, \[ |f(b)-f(a)|< M(b-a) \]

OpenStudy (hitaro9):

So |sqrt101-sqrt100| ends up being about .5, and (b-a) is 1 so you can throw htat out So ~.05 <M Why can M not be any number greater than .05?

OpenStudy (anonymous):

There is no bounds on M

OpenStudy (anonymous):

M can be one billion.

OpenStudy (hitaro9):

The online thing that I'm using said that a value of 1 for M was wrong.

OpenStudy (anonymous):

It only has a lower bounds: it can't be lower than the derivative.

OpenStudy (anonymous):

Well, the reason is that they want you to be more precise.

OpenStudy (anonymous):

M can be as big as you want, but you want M to be a low as possible in order to tighten the bounds around \(f(b)\) as much as possible.

OpenStudy (hitaro9):

So it's not that my answer was necesarily wrong, it was just that it wasn't good enough? Do I just need to guess at what they think is a small enough value?

OpenStudy (anonymous):

Well, there is a smallest value M can be.

OpenStudy (anonymous):

The smallest value M can be and still satisfy the requirements, is the the absolute extrema in the interval.

OpenStudy (anonymous):

They want the smallest M, most likely.

OpenStudy (hitaro9):

And that would be sqrt101 -sqrt 100 right? They don't allow me to enter in a squareroot, and that is an irrational number right?

OpenStudy (anonymous):

Nope, wrong. They want M to be the largest that \(f'(x)\) gets in the inverval \([a,b]\).

OpenStudy (anonymous):

Since \(f'(x)\) is an always decreasing function, then it's maximum would be at the beginning of the interval, in other words \(f'(a)\).

OpenStudy (anonymous):

Which evaluates to \(1/20\).

OpenStudy (hitaro9):

Oooohh

OpenStudy (hitaro9):

I see

OpenStudy (hitaro9):

That makes a lot of sense. Sorry I was being so dumb with this.

OpenStudy (hitaro9):

yeah. That was the right answer. Thank you so much dude